I went to burger king and bought a whopper with cheese for lunch the burger had 50g of fat ,45g of carbohydrates and 30g of protein what percentage of calories came from carbohydrate
Answer #1

There are 9, 4, 4 calories in 1 gram of fat, carbohydrates, protein respectively.

Calories obtained from fat of the food = 50× 9= 450 calories.

Calories obtained from carbohydrates of the food =45×4

=180 calories

Calories obtained from protein of the food = 30×4

= 120 calories

Total calories in the food = 450+180+120= 750 calories

Percentage of calories came from carbohydrates is

=180÷750×100 = 24 %